OpenStudy (anonymous):

Find the solutions to the system of equations. y=x^2+7x+7 y=x+2

OpenStudy (anonymous):

Can someone please work it out step by step?

OpenStudy (jdoe0001):

\(\bf \begin{array}{llll} {\color{red}{ y}}=x^2+7x+7\\ \bf {\color{red}{ y}}=x+2 \end{array}\implies x^2+7x+7=x+2\)

OpenStudy (anonymous):

There's your graph. The formula being graphed is: x^2 + 7x + 7 and x + 2 First intersection point is (-5, -3) Second intersection point is (-1, 1) It's hard to see on the graph but you can solve for it as well algebraically. Your 2 equations are: y = x^2 + 7x + 7 y = x + 2 Since both expressions on the right of the equal sign are equal to y, you can set them equal to each other to get: x^2 + 7x + 7 = x + 2 Subtract x and subtract 2 from both sides of the equation to get: x^2 + 7x - x + 7 - 2 = 0 Combine like terms to get: x^2 + 6x + 5 = 0 Factor this quadratic equation to get: (x + 5) + (x + 1) = 0 Solve for x to get: x = -5 x = -1 Substitute in either equation to solve for y. When x = -5, y = x + 2 becomes y = -3. When x = -1, y = x + 2 becomes y = 1. That gets your your intersection points of (-5, -3) and (-1, 1). You can also substitute in the other equation to get the same answer. Example: y = x^2 + 7x + 2 becomes y = (-5)^2 + 7(-5) + 7 which becomes y = 25 - 35 + 7 which becomes y = -10 + 7 which becomes y = -3.

OpenStudy (anonymous):

Thank you guys... ♥

OpenStudy (anonymous):

You're welcome again!! Xo!